Evaluate the following expressions:
log7 49
log2 1/32

ln 1/e^2
log10 0.001

Answer by KnightOwlTutor(293)

log7 49 =2 because 7^2=49
log2 1/32=-5 because 2^-5 =1/32
ln 1/e^2= -2 because e^-2
log10 0.001=-3 because 10^-3
The log is the exponent value